OpenStudy (calculusxy):

The components of vector [24, 7] are 24 and 7. Find the components of a vector that is three fifths as long as [24, 7].

OpenStudy (calculusxy):

@agent0smith

OpenStudy (agent0smith):

This is even easier than the last.

OpenStudy (calculusxy):

[14.4, 4.2]

OpenStudy (calculusxy):

I just wanted to make sure my answer :)

OpenStudy (agent0smith):

Find the length of the original vector Find the length of the new vector. Is the new one 3/5ths of the original?

OpenStudy (calculusxy):

I did 3/5 x 24/1 = 72/5 = 14.4 3/5 x 7/1 = 21/5 = 4.2

OpenStudy (agent0smith):

No i mean find the length using pythagorean theorem.